Solve 3x^4 −x^2 +8 = 0. Hint : Substitute x^2 = s

Thank you for helping me!!!
Answer #1

we have to Solve 3x4-x² + 8 = 0 Substituting x2 =s So, 324 x² +8=0 s+8=0 oo, 3.8² Soi SE -14)+ (-1)?-4.3.8 2.3 or S 1 7 1-36